Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 17     Entry time: 08 Mar 2019, 13:01
Author: Adriano Violante 
Location: General 
Severity: Information 
Subject: Minutes of the dispatch meeting 09.00 

SPB/SFX

  • Only minor issues being addressed with the support groups.

FXE

SQS

SCS

MID

  • Epix planning is advancing. List of activity being maintained and followed up by CY. Interface box for cables being modified by KS, CRs being implemented by AE. Ongoing. Goal date is 10.04.2019 to have the detector completely commissioned

HED

OPT Lasers

XRO

​​XPD

​​VAC

AE

  • for SQS there are issue with the Karabo/digitizers communication. Workaround exists, being fixed for good (AE + CAS).
  • Given the preliminary shutdown task list sent to readiness this week, it is clear that AE does not have the resources to cope witl all the requests. The tasks do not have priorities right now and this is a problem. In general, priorities and tasks that could be dropped are needed. Instruments, Support groups and PSPO cannot decide. PSPO will address the issue to HS/TT. 

ITDM

DET

  • Calibration pipeline stability test will be performed with SPB (during shifts). CAS/DET/SPB on it.

EET

CAS

  • Hotfix 2.3.6 (GUI server, data loggers bugs fixed) deployed in SA1,2, in SA3 not yet.

AoB

 

ELOG V3.1.4-7c3fd00